In an outbreak of salmonellosis, 25 persons from 25 different households became ill, in a community with a total population of 1000. Calculate the attack rate. Enter your response as a percent with one decimal place.
Answer #1

Attack rate = number of people ill ÷ number of people at risk

Number of ill people = 25

Number of people at risk = 1000

Attack rate = 25 ÷ 1000 = 0.025

To convert in percent multiply by 100

Attack rate = 0.025 × 100 = 2.5%